OpenStudy (jrmeyer):

A and B are constants such that (x+6)/x^2-3x = (A/x)+(B/x-3) . Find A-B.

jhonyy9 (jhonyy9):

so you need makeing the calcules by common denominator so for from the first fraction from denominator factoriz out x and will get ? can you calcule it ?

OpenStudy (jrmeyer):

Could you possibly solve it showing step by step?

jhonyy9 (jhonyy9):

so will be (x+6)/(x(x-3)) = (A/x ) +(B/(x-3) yes ?

OpenStudy (jrmeyer):

Would it be possible for you to solve the entire problem?

jhonyy9 (jhonyy9):

how do you think it ? i like help you understanding how you need solving the same exercises correct ok ?

jhonyy9 (jhonyy9):

so but you need cooperating what you know ok ?

OpenStudy (jrmeyer):

I'm helping someone out with this and I've been away from this type of math for years..that's why I was asking for a step by step, so I could explain it

jhonyy9 (jhonyy9):

ok so can you write here the needed calcules how you think it right ?

jhonyy9 (jhonyy9):

so than no in the first step (x+6) A B ------ = --- + ------ x(x-3) x x-3 this is ok ,yes ?

OpenStudy (jrmeyer):

yes

jhonyy9 (jhonyy9):

so than on the right side we need calcule the common denominator so what will be x(x-3) and for we can assuming these terms we need multiplie A by (x-3) and B by x and will get (x+6) (x-3)A + xB ------- = ----------- x(x-3) x(x-3) so there are on both sides the same ,equales common denominators than we can making the numerators equales the same and will get (x+6) =(x-3)A +xB make the calcule on the right side (x+6)=xA-3A +xB factoriz out the x and will get (x+6)=x(A+B) -3A assum to both sides 3A x+6+3A=x(A+B) divide both sides by x (x+6+3A) --------- = A+B x so but you need to find A-B how ?

jhonyy9 (jhonyy9):

opinion ?

jhonyy9 (jhonyy9):

so add to both sides -2B and will get it (x+6+3A) --------- -2B = A+B -2B x (x+6+3A-2B) ------------- = A -B x ok ?

OpenStudy (jrmeyer):

Great - thanks for all your help...greatly appreciated!

jhonyy9 (jhonyy9):

was my pleasure yw so you have thougt it the same ?